OpenStudy (gabylovesyou):

The formula for degrees Celsius is: C = 5/9(F − 32), where F stands for degrees Fahrenheit. Part 1: Solve the equation for F and show all steps. Part 2: Determine how many degrees Fahrenheit 15 degrees Celsius is.

OpenStudy (anonymous):

plug it in :)

OpenStudy (anonymous):

\[C=\frac{5}{9}(F-32)\] \[\frac{9}{5}C=F-32\]

OpenStudy (anonymous):

\[\frac{9}{5}C+32=F\]

OpenStudy (anonymous):

this is the formula you use... °C x 9/5 + 32 = °F swap C for 12... 12 x 9/5 + 32 = °F then multiply 12 and 9/5... 108/5 + 32=*F make 108/5 to a mixed number... 21 and 3/5 + 32= *F then add... 53 3/5 *F

OpenStudy (anonymous):